what is the formula for probability of anumber out of many numbers,example probability of 6 numbers out of 42 numbers
Answer by stanbon(75887)

The probability of an event
= (# or ways to succeed)/(# of ways to succeed + # of ways to fail)
-----------------------------------------------------------------------
Your Problem:
If you have a bag with 6 red marbles and a total of 42 marbles,
the probability of selecting a red marble from the bag is
6/42 = 3/21 = 1/7
=========================
Cheers,
Stan H.